# PrisonBreak: Jailbreaking Large Language Models with Fewer Than Twenty-Five Targeted Bit-flips

### v1

@@ 7.3 Comparison of Harmful Responses
In this section, we conduct a qualitative analysis of our attack’s outputs and how they differ from the original model and baseline methods. Table 4 provides examples that highlight the general trends we observe for each method.
Our attack generates higher-fidelity responses than Baseline BFA. Across all models, PrisonBreak induces a jailbroken state largely avoidant of the failure modes presented in §5.1. Models rarely backtrack affirmative responses, and although we find a few cases of overfitting, they are still accompanied by harmful content in most cases. Conversely, Baseline BFA is severely susceptible to overfitting and commonly yields no harmful content but rather the repetition of a few phrases. Despite these undesirable artifacts, the HarmBench evaluator tends to still report these instances as harmful, likely because the model does not explicitly deny the query and repeatedly affirms it. As a result, the ASRs reported for Baseline BFA likely overestimate the amount of harmful content produced.
Comparison with non-BFA methods. While all three attacks (excluding Baseline BFA) often generate high-fidelity content, they differ with respect to output style. We find that GCG often involves portions of the adversarial suffix, for example, it formats the example response in “Irish colloquial style” (the suffix contains the word “Irish”). While this does not impede output quality, it sometimes overvalues the suffix and ignores the harmful request, possibly playing a role in the prompt-based method’s lower ASR. The responses from PrisonBreak are semantically closest to Ortho. Table 4 highlights how model’s attacked by these methods provide very similar steps in the step-by-step instructions, and generally respond to queries with comparable ideas. We interestingly observe that responses from Ortho tend to be more descriptive and contain more information. This is likely because our proxy dataset lacks this verbosity, instead utilizing more direct answers. However, PrisonBreak still produces substantially harmful responses, and we leave the evaluation of different proxy datasets as future work.
@@ 7.4 Impact of Attack Configurations
We assess how sensitive our attack is to different configuration choices. Figure 2 shows our results on Llama2-7B. Additional results are shown in Appendix D.
@@
Impact of the proxy dataset size. Our default proxy dataset size is 41—a small number compared to prior work which uses 50–2500 samples to compute the attack objective. The leftmost figure in Figure 2 shows the ASR and Acc. for different proxy dataset sizes in jailbreaking Llama2-7B. We observe similar output quality across all dataset sizes and find that any dataset size over 5 achieves an ASR above 85%, indicating that a large dataset is not necessary for successful jailbreaks. However, we find that these smaller dataset sizes are often dependent on the specific samples they contain. We use the full proxy dataset in our evaluation, to promote consistent results between models.
Figure 2: The impact of proxy dataset size, number of critical weights, utility score threshold, and number of bit-flips on ASR and Acc. in Llama2-7B.
@@
Impact of $k$, the number of critical weights selected per layer during the search process. Using a smaller $k$ reduces the total computations, but may compromise attack success and model utility. The second figure from the left shows the impact of $k$ on ASR and Acc. As $k$ increases from 10 to 150, ASR increases from 73.6% to 95.0%. But, Acc. decreases marginally from 51.8% to 50.7%, and the computation required for our attack increases proportionally. Despite the reduction in accuracy, the quality of harmful responses remains similar across $k$. Our preliminary investigation finds that $k=$ 50–100 balances the time it takes to search target bits while increasing ASR only with a marginal reduction of Acc. We set $k=100$ in our main attack results.
@@
Impact of $\tau$, the utility score threshold in selecting target bits. We vary $\tau$ in $\{0.001,0.01,0.1,1\}$, with smaller values indicate that flipping the selected bit results in only a slight decrease in model utility. The third figure from the left shows the ASR and Acc. across these values. It shows that $\tau$ does not affect ASR, but it reduces accuracy from 52.9–50.3% as $\tau$ increases. Manual analysis of generations across $\tau$ values shows that a smaller threshold produces higher-fidelity responses for both benign and harmful queries. For example, the generations of Llama2-7b with $\tau\!=\!1$ contain much more overfitting than those with lower thresholds, even at $\tau\!=\!0.01$. In our attacks, we set $\tau=0.01$.
@@
Impact of the number of target bits. §7.2 shows our attack to be effective for at most 25 bit-flips, but an attacker may also choose to flip fewer bits. We explore the impact of 0–25 bit flips on both ASR and Acc. The rightmost figure shows that ASR increases substantially first from 1.3–62.9% as we increase target bits from 0–10, and then gradually to 86.8% after all 25. The sudden increase in ASR at six bit-flips indicates that cumulative bit flipping has sufficiently altered the model’s output distribution toward the jailbreak objective. ASR saturates after 20 bit-flips, with flips 20–25 increasing ASR by only 5% points. For Acc., we observe a slight decrease from 52.2–50.3% as more bits are flipped. Given the marginal change (less than 2%), the impact on model behavior appears minimal. This non-proportional trade-off between ASR and Acc. enables the attacker to achieve both jailbreak success and stealthiness.
Models # of Bit-Flips High (RV $>$ 10 ppm) Mid (1 ppm $<$ RV $\leq$ 10 ppm) Low (RV $\leq$ 1 ppm) A17 (65) A8 (43) A12 (20) A19 (5.4) D1 (3.1) A11 (1.5) A3 (0.8) C4 (0.4) B2 (0.01) Vicuna-7B-v1.3 5 100% 100% 100% 99.6% 96.3% 38.1% 10.7% 0.7% 0% Vicuna-7B-v1.5 12 100% 100% 100% 93.5% 59.8% 0% 0% 0% 0% Vicuna-13B-v1.5 20 100% 100% 100% 97% 76% 0.4% 0% 0% 0% Llama2-7B 25 100% 100% 100% 19.4% 0.4% 0% 0% 0% 0% Llama2-13B 18 100% 100% 100% 96.9% 74.8% 0.1% 0% 0% 0% Llama3-8B 20 100% 100% 100% 94.7% 63.2% 0.1% 0% 0% 0% Qwen2-1.5B 25 100% 100% 100% 93.8% 53.5% 0% 0% 0% 0% Qwen2-7B 25 100% 100% 100% 93.5% 56.1% 0% 0% 0% 0%
@@
Table 5: PrecisionHammer Results on DDR4 Modules. We show the Rowhammer vulnerability (RV) (ppm) of each module in parentheses. For each RV level, we show the results for three representative modules. Results for all modules are presented in Table 10.
8 Rowhammer Exploitation
This section demonstrates the practical exploitation of PrecisionHammer. Rather than running the online procedure on a few sets of commodity hardware [1, 7], we follow the prior work [2] and use simulations to assess the exploitability across various DRAM modules with different levels of Rowhammer vulnerability (RV).
Experimental setup. We conduct Monte Carlo simulations based on Rowhammer profiles from prior work [22]. Specifically, we evaluate 40 DDR4 and 16 LPDDR4X DRAM modules with different RVs, defined as the ratio of cells susceptible to bit-flip to the total number of DRAM cells. We categorize 40 DDR4 modules into three levels of RVs:
@@ 8 Rowhammer Exploitation
High-RV modules have vulnerability levels comparable to those used in prior studies [1, 14, 11], while Mid- and Low-RV modules have 3-14$\times$ and 18-2560$\times$ fewer vulnerable cells than the modules evaluated in prior attacks. All 16 LPDDR4X modules have high RVs. We evaluate the system in a single-channel configuration for both DDR4 and LPDDR4X modules. Detailed descriptions of these modules are available in prior work [22].
Models # of Bit-Flips 15/16 Modules (RV $\geq$ 19.5) C5 (10.8) Vicuna-7B-v1.3 5 100% 100% Vicuna-7B-v1.5 12 100% 100% Vicuna-13B-v1.5 20 100% 98% Llama2-7B 25 100% 80% Llama2-13B 18 100% 100% Llama3-8B 20 100% 100% Qwen2-1.5B 25 100% 100% Qwen2-7B 25 100% 100%
Table 6: PrecisionHammer Results on LPDDR4 Modules.
[…]
@@
Hardware-agnostic attacker. We first evaluate the scenario in which the attacker identifies a set of target bits and aims to flip them across multiple DRAM modules. Table 5 summarizes the attack success for three representative DDR4 modules at each RV level. Please refer to Table 10 in the Appendix for our full results. Our hardware-agnostic attack is consistently successful on high-RV DRAMs, in alignment with prior work [1, 14, 11, 7], especially as our attack requires a similar or fewer number of bit-flips. The attacks are reasonably successful on Mid-RV DRAMs, demonstrating a success rate of 96.3% for Vicuna-7B-v1.3 on D1, a module with 5.4$\times$ less RV than those used in prior work. Even for Qwen2-7B, which requires the full 25 bit-flips, the attack achieves a success rate of 56.1%. But, we are less successful for Llama2-7B against Mid-RV modules as many target bits share the same offsets, requiring multiple Rowhammer cells with identical offsets within a DRAM row. Given the uniform distribution of Rowhammer cells, this lowers the probability of finding exploitable bits for all 25 targets. The attacks are not successful on Low-RV DRAMs, except for Vicuna-7B-v1.3, as these modules are significantly more resilient to Rowhammer with far fewer exploitable bits (a few hundred to several thousand).
@@
We also evaluate our attack on low-power (LP) DDR4X modules. In Table 6, our hardware-agnostic attacks are consistently successful. This is because LP devices tend to be more vulnerable to Rowhammer due to their power and area constraints [22, 42]. The only exception is the C5 module, with the lowest attack success of 80% (Llama2-7b), as it has the lowest RV.
@@
Table 7: Hardware-aware search results. ASR of PrisonBreak on hardware profiles A11 and C4, using our hardware-aware search. The Rowhammer vulnerability (ppm) of each module is in parentheses.
Hardware-aware attacker. We now evaluate the effectiveness of hardware-aware attacks against secure DRAM modules with only a few thousand vulnerable cells. We consider the A11 (Mid-RV) and C4 (Low-RV) modules, where the hardware-agnostic attack achieved substantially low simulation success (between 0–38.1% across all models). Table 7 shows the ASR achieved by the hardware-aware attacker, computed on HarmBench as in §7. We do not report the simulation success as these attacks produce bit-flip sequences that are always flippable in the target modules. The results show that the hardware-aware attacker achieves a high ASR across both modules, comparable to the upper-bound results we obtain in §7. Even in this constrained setting, PrisonBreak can still jailbreak models with 25 or fewer bit-flips. For reference, the A11 and C4 modules are 12–46$\times$ more secure than those used in prior work [1].
@@ 9 Understanding the Vulnerability
We first analyze the model structure to understand which components contribute to achieving the jailbreak. We next conduct a behavioral analysis to identify the internal mechanisms of a model that our attack exploits.
@@ 9.1 Structural Analysis
To assess the vulnerability, we analyze the bit-flips made during our attack that lead to the highest reduction in our jailbreak objective. These bits are the most likely to contribute to successful jailbreaks. For a representative analysis, we examine the top 1% of bits identified in each iteration of BitFinder, ranked via the jailbreak score defined in Def. 5.1. We consider all 16 bit locations in each critical weight. Figure 3 presents the proportion of target bits distributed across Llama2-7B components.
@@
Figure 3: Structural analysis results. The proportion of target bits across bit locations, layer types, and layer positions of Llama2-7b. We define target bits as the top 1% evaluated by BitFinder and report the proportions across 25 iterations.
@@
Bit locations. The leftmost figure confirms our insight that large parameter changes, specifically induced by flipping the three most significant bits in the exponent, lead to the greatest reduction in our jailbreaking objective (see §5.3). Within the top 1% bits across each bit location in the half-precision floating-point weight value, we see that the three most significant exponent bits makeup over 94% of all top bits flips across iterations. Of these three bit locations, the most significant bit (MSB) in the exponent is likely chosen (with a proportion of 48.0%) in the first round, while the second bit is more likely for the following iterations with proportions of 42.4–60.4%. This is because, while flipping the second bit induces smaller parameter changes than the first, it prevents excessive changes that could often reduce a model’s generation capabilities significantly. Our findings corroborate the actual bits selected by BitFinder: 31.5% are MSBs, and 64% are the second exponent bits.
@@
Types of layers composing the Transformer block. Given that diverse functions of layers compose the Transformer block—a core architectural component of modern language models, BitFinder identifies more candidate bits in certain layers than others. The middle figure shows the proportion of the top 1% bits found in each layer type. Overall, the linear layer for the value in multi-head attention (attn.v) has noticeably higher proportions, between 25.2–35.0%. Like the bit location analysis, this indicates the layers where bits are actually flipped to jailbreak the model, as 28.5% of all bits selected in §7 come from these layers. A possible explanation for this discrepancy is that PrisonBreak primarily affects how a model responds to harmful queries (i.e., not refusing to answer) rather than how it recognizes these queries. In multi-head attention, “how to respond” corresponds to the value layers, while “how to recognize” aligns with the key and query layers.
@@
Layer positions. Prior work [63] empirically shows that in Transformer-based models, earlier layers handle basic concepts like linear word order, while syntactic information is most prevalent in the middle layers, and final layers are task-specific. Here we also study the relationship between this observation and the likelihood of selecting target bits across different layer positions. We define three categories for Llama2-7B: earlier layers that are in transformer blocks 1–10, middle layers in blocks 11–22, and final layers in blocks 23–32. The rightmost figure shows the proportions of the top 1% bits in these categories. The earlier layers are less likely chosen, accounting for less than 25% of top 1% bits in most iterations. The latter two categories account for the majority of the top 1% bits, particularly the middle layers, with proportions ranging from 33.5% to 63.1%. Our findings corroborate the prior work’s observations: in early iterations, BitFinder is more likely to select target bits from the final layers, as the attack focuses on a task of generating responses to harmful queries. But in the subsequent iterations, the attack shifts to choosing target bits from the middle layers, focusing on generating high-quality responses while maintaining attack success (the attack objective).

### v2

@@ 6.3 Comparison of Harmful Responses
In this section, we conduct a qualitative analysis of our attack’s outputs and how they differ from the original model and baseline methods. Table IV provides examples that highlight the general trends we observe for each method.
Our attack generates higher-fidelity responses than Baseline BFA. PrisonBreak induces a jailbroken state largely avoidant of the failure modes presented in §5.1. Models rarely backtrack affirmative responses, and although we find a few cases of overfitting, they are still accompanied by harmful content. Conversely, Baseline BFA is severely susceptible to overfitting and commonly yields no harmful content but rather the repetition of a few phrases. Despite these undesirable artifacts, the HarmBench evaluator tends to still report these instances as harmful, likely because the model does not explicitly deny the query and repeatedly affirms it. As a result, the ASRs reported for Baseline BFA likely overestimate the amount of harmful content produced.
Comparison with non-BFA methods. While all three attacks (excluding Baseline BFA) often generate high-fidelity content, they differ with respect to output style. GCG frequently involves portions of the adversarial suffix, e.g., the model formats the example response in “Irish colloquial style” (the suffix contains “Irish”). Models sometimes overvalue the suffix and ignore the harmful request completely, possibly playing a role in the prompt-based method’s lower ASR. The responses from PrisonBreak are semantically closest to Ortho; models attacked by these methods provide very similar step-by-step instructions and generally respond to queries with comparable ideas.
6.4 Exploitability Under Black-Box Settings
We also evaluate the effectiveness in black-box scenarios where an adversary may not have white-box access to the exact model they wish to jailbreak. Here, the attacker can exploit the standard paradigm of LLM development: pretraining-then-finetuning. Because pre-training is the primary computational bottleneck, many practitioners open-source pre-trained LLMs for others to fine-tune [30, 25, 29, 59, 31, 60]. This practice has led to a proliferation of models that share components, such as architectures, pre-training, fine-tuning, or post-training alignment procedures. As a result, an attacker with white-box access to an open-source model (source) can identify effective bit-flip locations in memory and transfer them to a target model in a black-box setting without any access to its parameters.
[…]
@@ 6.5 Impact of Attack Configurations
We analyze how sensitive our ASR is to different attack configurations. We show our results on Llama2-7B in Figure 2. Please refer to Appendix B for another model.
@@
Impact of the proxy dataset size. Our default proxy dataset size is 41—a small number compared to prior work which uses 50–2500 samples to compute the attack objective. The leftmost figure shows the ASR and Acc. for different proxy dataset sizes in jailbreaking Llama2-7B. We observe similar output quality across all dataset sizes and find that any dataset size over 5 achieves an ASR above 85%, indicating that a large dataset is not necessary for successful jailbreaks. However, we find that smaller datasets are often dependent on the specific samples they contain; we use the full proxy dataset to promote consistent results between models.
@@
Impact of $k$, the number of critical weights selected per layer during the search process. A smaller $k$ reduces the total computations, but may compromise attack success and model utility. The second figure from the left shows the impact of $k$ on ASR and Acc. As $k$ increases from 10 to 150, ASR increases from 73.6% to 95.0%. But, Acc. decreases marginally from 51.8% to 50.7%, and the computation required for our attack increases proportionally. Despite the reduction in Acc., the quality of harmful responses remains similar across $k$. Our preliminary investigation finds that $k=$ 50–100 balances search time while increasing ASR only with a marginal reduction of Acc. We use $k=100$.
@@
Impact of $\tau$, the utility score threshold used when selecting target bits. We vary $\tau$ in $\{0.001,0.01,0.1,1\}$, with smaller values indicating that flipping the selected bit only slightly decreases model utility. The third figure from the left shows the results: $\tau$ does not affect ASR, but as it is increased Acc. is reduced from 52.9% to 50.3%. Manual analysis of outputs across $\tau$ values shows that a smaller threshold produces higher-fidelity responses for both benign and harmful queries. For example, the generations of Llama2-7b with $\tau\!=\!1$ contain much more overfitting than those with lower thresholds, even at $\tau\!=\!0.01$. In our attacks, we set $\tau=0.01$.
@@
Impact of the number of target bits. §6.2 shows our attack to be effective for up to 25 bit-flips, but an attacker may also choose to flip fewer bits. We explore the impact of 0–25 bit flips on ASR and Acc. The rightmost figure shows that ASR increases first from 1.3$\rightarrow$62.9as we increase target bits from 0$\rightarrow$10, and then gradually to 86.8% after all 25. The sudden increase in ASR at six bit-flips indicates that cumulative bit flipping has sufficiently altered the model’s output distribution toward the jailbreak objective. ASR saturates after 20 bit-flips, with flips 20–25 increasing it by only 5% points. For Acc., we observe a decrease from 52.2$\rightarrow$50.3as more bits are flipped. Given the marginal change (less than 2%), the impact on model behavior appears minimal.
7 Rowhammer Exploitability
We now demonstrate the Rowhammer exploitability of PrisonBreak. Rowhammer exploitation typically breaks down into three critical phases: (1) templating, (2) memory massaging, and (3) reflipping. Templating identifies DRAM rows and bit positions that reliably flip under repeated access (hammering). Attackers test different aggressor rows using various access patterns to discover potentially exploitable bit positions [39]. Prior studies show that templating takes anywhere from several hours to days [1, 39] but is performed offline and only once per machine. Once exploitable bit positions are identified, memory massaging relocates the pages containing the target bits in a model into physical DRAM rows containing these vulnerable bits. This is achieved through direct memory control and hammering [36], or by repeatedly allocating and freeing “dummy” pages to force the target into the desired location [61, 62]. With the target page in place, the attacker reflips the vulnerable bit to corrupt model parameters. This final phase is performed online and typically completes in a few minutes [1, 63, 12]. Attacks are carefully crafted to hammer only templated aggressors [62, 64, 65], so as to avoid extraneous flips that could crash the process or degrade model utility.
Prior work has shown that precise Rowhammer exploitation is feasible for neural networks [1, 12, 7, 11], even when flipping up to 59$\times$ more bits [11] than PrisonBreak requires. Rather than re-demonstrating Rowhammer on limited commodity hardware, we follow recent work [66] and use simulations to assess the exploitability across dozens of DRAM modules. This prevents our results from being an one-time exploitation and instead allows us to comprehensively assess generalizability to many DRAM modules, in line with prior large-scale Rowhammer studies [39].
@@ 7 Rowhammer Exploitability
TABLE VI: Rowhammer results on DDR4 and LPDDR4 modules. We show the Rowhammer vulnerability (RV) (ppm) of each module in parentheses. For each RV level, we report the RPS for representative modules. Results for all DDR4 modules are in Table X.
Models # of Bit-Flips DDR4 LPDDR4 High-RV Mid-RV High-RV A17 (65) A8 (43) A12 (20) A19 (5.4) D1 (3.1) A11 (1.5) 15/16 Modules ($\geq$19.5) C5 (10.8) Vicuna-7B-v1.3 5 100% 100% 100% 99.6% 96.3% 38.1% 100% 100% Vicuna-7B-v1.5 12 100% 100% 100% 93.5% 59.8% 0% 100% 100% Vicuna-13B-v1.5 20 100% 100% 100% 97% 76% 0.4% 100% 98% Llama2-7B 25 100% 100% 100% 19.4% 0.4% 0% 100% 80% Llama2-13B 18 100% 100% 100% 96.9% 74.8% 0.1% 100% 100% Llama3-8B 20 100% 100% 100% 94.7% 63.2% 0.1% 100% 100% Llama3-70B 18 100% 100% 100% 95.1% 68.2% 0.1% 100% 100% Tulu3-8B 13 100% 100% 100% 95.4% 66.3% 0% 100% 99% Qwen2-1.5B 25 100% 100% 100% 93.8% 53.5% 0% 100% 100% Qwen2-7B 25 100% 100% 100% 93.5% 56.1% 0% 100% 100%
7.1 Experimental Setup and Methodology
To test the exploitability, we assume the previously demonstrated templating, massaging, and re-flipping capabilities [1, 7, 12] and simulate an attacker’s ability to locate and re-flip vulnerable bit locations identified by PrisonBreak. Specifically, we conduct Monte Carlo simulations based on Rowhammer profiles from Blacksmith [39]. We evaluate 40 DDR4 and 16 LPDDR4X (Low-Power DDR4 eXtension) DRAM modules from four manufacturers. Each DRAM module consists of a profile detailing the number of cells vulnerable to Rowhammer and the number of bit-flips possible in each flip direction (0$\rightarrow$1 or 1$\rightarrow$0). The total number of vulnerable cells determines a module’s Rowhammer vulnerability (RV), defined as the ratio of cells susceptible to bit-flips to the total number of DRAM cells. We categorize the 40 DDR4 modules into three RV levels:
@@ 7.1 Experimental Setup and Methodology
High-RV modules show vulnerability levels comparable to those exploited in prior BFA studies [1, 12, 11]; Mid- and Low-RV modules have 3–14$\times$ and 18–2560$\times$ fewer vulnerable cells than the High-RV modules. All 16 LPDDR4X modules have High-RVs. Our simulation considers a system with a single-channel configuration for all modules. Descriptions of these modules are available in prior work [39].
Methodology. In each iteration of the Monte Carlo simulation, we randomly initialize the Rowhammer cells based on the RV of each DRAM module. Cell locations are sampled uniformly, because prior work [1] found that flippable bit-flip locations are uniformly distributed, and the Rowhammer profiles we reference [39] do not specify the exact locations of vulnerable cells. After the cell initialization, we check if there are any exploitable Rowhammer cells for our target bits, following a process similar to the Hammertime framework [63]. Each target bit must be matched to a cell with the same offset and flip direction, otherwise it is not considered flippable. This constraint mimics the attacker’s memory massaging step, where they can only massage target bits into specific DRAM cells. The attack is successful if we find exploitable locations for all required target bits necessary to jailbreak the LLM, modeling a successful memory massaging attempt by the attacker that can be followed by Rowhammer. We run the simulations for 100,000 iterations.
Metrics. We employ two metrics to quantify the Rowhammer exploitation. The first is the Rowhammer success rate (RSR), estimated using our Monte Carlo simulation. RSR is defined as the ratio of successful trials to the total number of trials. This metric applies to the hardware-agnostic attacker, who does not profile the DRAM before attacking, making their RSR dependent on the RH vulnerability of the DRAM module. The second metric is the attack success rate (ASR) introduced in §6.1. We use this metric for the hardware-aware attacker; this attacker always selects flippable target bits, making Monte Carlo simulations unnecessary.
7.2 Evaluation Results
@@
Hardware-agnostic attacker. We first evaluate the scenario where the attacker identifies a set of target bits and aims to flip them across multiple DRAM modules. We use the same target bits as in §6.2. The third and fourth columns in Table VI show the RSR for three representative DDR4 modules at the High- and Mid-RV levels; we omit the Low-RV results due to low RSR (0010101010%) and present them with the full results in Table X in the Appendix. Our hardware-agnostic attack achieves an RSR of 100% across all High-RV DRAMs, in alignment with prior work [1, 12, 11, 7], especially as our attack requires a similar or fewer number of bit-flips. The attack is reasonably successful on Mid-RV DRAMs, with an RSR of 96.3% for Vicuna-7B-v1.3 on D1, a module with 5.4$\times$ less RV than those used in prior work. Even for Qwen2-7B, which requires 25 bit-flips, we achieve an RSR of 56.1%. But, we are less successful for Llama2-7B against Mid-RV modules as many target bits share the same offsets, requiring multiple Rowhammer cells with identical offsets within a DRAM row. Given the uniform distribution of vulnerable Rowhammer cells, this lowers the probability of finding exploitable locations. The attacks are not successful on Low-RV DRAMs, as these modules are significantly more resilient to Rowhammer with far fewer exploitable bits (at most several thousand).
@@
The rightmost column in Table VI shows our results on LPDDR4X modules. Our hardware-agnostic attack is consistently successful with an RSR of 80–100% across all settings. This is because LP devices are more vulnerable to Rowhammer due to power and area constraints [39, 37].
TABLE VII: Hardware-aware attack results. We report the ASR of PrisonBreak on hardware profiles A11 and C4.
@@
Hardware-aware attacker. This attacker identifies target bits that map to vulnerable cells and can succeed even against secure DRAM modules. We evaluate against the A11 (Mid-RV) and C4 (Low-RV) modules, where the hardware-agnostic attack achieved very low RSR (0–38.1%). Table VII shows the ASR for the hardware-aware attacker, computed on HarmBench as in §6. We do not report the RSR as this attack produces bit-flip sequences that are always flippable in the target DRAM. The hardware-aware attacker achieves a high ASR on both modules, comparable to the upper bound results we obtain in §6. Even in constrained settings, PrisonBreak can still jailbreak models with up to 25 bit-flips. For reference, the A11 and C4 modules are 12–46$\times$ more secure than those used in prior work [1].
@@ 8 Understanding the Vulnerability
We now analyze the model structure to identify components enabling the jailbreak and perform a behavioral analysis to find internal mechanisms exploited by our attack.
@@ 8.1 Structural Analysis
Here we analyze bit-flips that led to the highest reduction in our jailbreak objective,
particularly the top 1% of bits ranked by the jailbreak score. We consider all 16 bit locations in each critical weight. Figure 3 shows the proportion of target bits across different components of Llama2-7B.
@@
Figure 3: Structural analysis results. The proportion of target bits across bit locations, layer types, and layer positions of Llama2-7b. We consider the top 1% of target bits evaluated by our offline procedure and report the proportions across 25 iterations.
Figure 4: Behavorial analysis results. For each attack: the cosine similarity between HarmBench activations and the refusal direction across all layers of Vicuna-13b-v1.5 (left) and the UMAP visualization of activations from Alpaca (middle) and HarmBench (right) from the last layer of Qwen2-7b.
@@
Bit locations. The leftmost figure confirms that large parameter changes, induced by flipping the three most significant bits (MSBs) in the exponent, lead to the greatest reduction in our objective. Within the top 1% bits across each bit location in the half-precision floating-point weight value, we see that the three most significant exponent bits makeup over 94% of all top bits flips across iterations. Of these three bit locations, the MSB in the exponent is likely chosen (with a proportion of 48.0%) in the first round, while the second bit is more likely for the following iterations with proportions of 42.4–60.4%. This is because, while flipping the second bit induces smaller parameter changes than the first, it prevents excessive changes that often reduce a model’s generation capabilities. Our findings corroborate the actual bits selected: 31.5% are MSBs, and 64% are the second exponent bits.
@@
Types of layers composing the Transformer block. Our attack identifies more candidate bits in certain Transformer layers than others. The middle figure shows the proportion of the top 1% bits found in each layer type. Overall, the linear layer for the value in multi-head attention (attn.v) has relatively higher proportions (25.2–35.0%); notably, 28.5% of all bits selected in §6 come from these layers. This may be because PrisonBreak primarily affects how the model responds to harmful queries (i.e., not refusing to answer) rather than how it recognizes them. In multi-head attention, “how to respond” corresponds to the value layers, while “how to recognize” aligns with the key and query layers.
@@
Layer positions. Prior work [67] shows that in Transformer-based models, earlier layers handle basic concepts like word order, middle layers focus on syntax, and final layers are task-specific. Building on this, we analyze the relationship between layer positions and the likelihood of selecting target bits in Llama2-7B. We categorize layers into early (blocks 1–10), middle (blocks 11–22), and final (blocks 23–32). The rightmost figure shows the results. Early layers are chosen less often, accounting for under 25% of top bits in most iterations. Middle and final layers dominate, particularly with middle layers contributing 33.5–63.1%. Our findings corroborate the prior work: in early iterations, the attack is more likely to select bits from the final layers, as it focuses on the task of generating harmful responses. In the later iterations, the attack shifts to choosing target bits from the middle layers, focusing on generating high-quality responses while maintaining attack success (the attack objective).